A box contains 3 red balls and 5 white balls. One ball is chosen at random from the box and is not returned to the box. A second ball is now chosen at random from the box.
- (i)Find the probability that both balls chosen are red.[1]
- (ii)Show that the probability that the balls chosen are of different colours is 15/28.[2]
- (iii)Given that the second ball chosen is red, find the probability that the first ball chosen is red.[2]
- (iv)The random variable X denotes the number of red balls chosen. Draw up the probability distribution table for X.[2]
- (v)Find Var(X).[3]
Mark scheme
Marking points from the official mark scheme (Oct/Nov 2019). The mark or mark type is shown in brackets.
- [B1] 3/8 × 2/7 = 3/28 OE
- [M1] Method shown, numerical calculations identified, may include replacements
- [A1] AG, Fully correct calculations
- [M1] Method shown, numerical calculations identified, may include replacements
- [A1] AG, Fully correct calculations
- [M1] Conditional probability formula used consistent with their probabilities or correct
- [A1] OE
- [B1] Probability distribution table with correct values of x and at least one correct probability placed. Extra x values allowed with probability of zero stated.
- [B1FT] Fully correct FT P(2) = their (i), P(1) = their (ii), Σ(p) = 1.
- [B1] May be implied by use in variance formula
- [M1] Substitute into correct variance formula, must have ' - their mean²' Must be for 2 or more non-zero x-values
- [A1] Correct final answer
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →